OpenStudy (anonymous):

Which equation describes the function shown on the graph? y = -3x + 4 y = 3x – 4 y = 1 3 x – 4

OpenStudy (compassionate):

There is no graph.

OpenStudy (precal):

can you tell me the value of your y intercept

OpenStudy (precal):

ie can you tell me where your line crosses the y axis?

OpenStudy (precal):

ok let me know when you would like my help

OpenStudy (anonymous):

help?

OpenStudy (precal):

still waiting for you to answer my question, this site is called openstudy not instant answers

OpenStudy (anonymous):

srry was trying to figure out on own when you said that... 4

OpenStudy (anonymous):

actually -4 not 4

OpenStudy (precal):

yes it crosses at -4 this is called your y-intercept y=mx+b so far you know the solution is y=mx-4 because your b value is -4 remember we call b your y intercept

OpenStudy (precal):

so let's figure out the slope

OpenStudy (precal):

that is your m value

OpenStudy (precal):

your slope is either a positive slope, negative slope, zero slope or undefined slope read you graph from left to right and tell me if it is going up (positive slope) down (negative slope) horizontal (0 slope) vertical (undefined slope)

OpenStudy (precal):

ok so is the line going up or down?

OpenStudy (anonymous):

up

OpenStudy (precal):

ok that means your number in front of x is positive

OpenStudy (precal):

do you know how to calculate the slope using rise over run?

OpenStudy (anonymous):

ya... learned that in 5th grade

OpenStudy (precal):

ok pick 2 points and tell me what your slope is

OpenStudy (anonymous):

run: 3 rise:1 (-6,-6) to (0,-4) its 1/3=.66 repeting

OpenStudy (precal):

you got it y=(1/3)x-4